Convert (5B5)12 = (          )10
Answer #1

Solution :

The given value is : (5B5)12=(?)10

Now the base 10 value can be given by ,

based on the position of the digit, The digit is multiplied by 12 power the position value ,starting from right (the first position value from right is 0)

=> (5*122)+(B*121)+(5*120)

=(5*144)+(11*12)+(5*1)

=(720)+(132)+(5)

=720+132+5

=857

So 857 is the required base 10 value

it implies,

(5B5)12=(857)10

If you find my answer helpful please give thumbs up,Thank you.